OpenStudy (anonymous):

Given the linear equation y=1/2x-5, find the y-coordinates of the points (-8),(-2) and (4).Please show all your work.Plot these points and graph the linear equation.Submit your graph.

OpenStudy (aroub):

y=1/2x-5 (-8, ) , (-2, ) and ( 4, ) How do you find the y-coordinates? By susbstituting the x- coordinates in the equation. Like this: lets try to find the first point y=1/2x-5 y=1/2(-8)-5 y=-4-5 y=-9 We found the first point which is (-8,-9) Can you do the rest now?

OpenStudy (anonymous):

No

OpenStudy (aroub):

Why?

OpenStudy (anonymous):

I am mad dumb.

OpenStudy (aroub):

Why don't you ATLEAST try?

OpenStudy (anonymous):

I did and I get something else somehow